explain application of gauss’s law and electric potential energy
 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Gauss Lawstates that the total electric flux out of a closed surface is equal to the charge enclosed divided by the permittivity. The electric flux in an area isdefinedas the electric field multiplied by the area of the surface projected in a plane and perpendicular to the field.

Application:
The intensity of the electric field near a plane sheet of charge is E = σ/2ε0K where σ = surface charge density.

Theintensity of the electric fieldnear a plane charged conductor E = σ/Kε0in a medium of dielectric constant K. If the dielectric medium is air, then Eair= σ/ε0.

The field between two parallel plates of a condenser is E =σ/ε0, whereσ is the surface charge density.

Thanks